OpenStudy (anonymous):

given that P(A) = .4, P(B) = .8, and P(A and B) = .40; determine P(A/B) ***the / is a straight line in the problem, not slanted

OpenStudy (a_clan):

p(A|B) = p(A and B)/p(B)

OpenStudy (anonymous):

so the answer is .5?

OpenStudy (a_clan):

yes

OpenStudy (anonymous):

yay! thank u so much! can u help me with another one?

OpenStudy (a_clan):

will try

OpenStudy (anonymous):

ok thanks, it is... if P(A or B) = .06, P(A) = .03, and P(B) = .07; determine P(A and B)

OpenStudy (a_clan):

P(A or B) = P(A) + P(B) - P(A and B) So P(A and B) = P(A) + P(B) -P(A or B)

OpenStudy (anonymous):

so .04?

OpenStudy (a_clan):

right

OpenStudy (anonymous):

cool, thanks a bunch